OpenStudy (anonymous):

please I need a slution for this : y"-4y'+3y=exp(x)+exp(-x)

OpenStudy (amistre64):

what do you get for the homogenous part?

OpenStudy (anonymous):

A*exp(4x) + B*exp(x) what I need is the particular solution

OpenStudy (amistre64):

your roots are 3 and 1, but ill assume thats a typo there is a trial and error method; assume a y = exp(x) + b exp(-x) is a solution find y' and y'' of those

OpenStudy (amistre64):

shoulda been an a in front of the first one

OpenStudy (anonymous):

yes yes excuse me it's 3 and 1

OpenStudy (amistre64):

y = a exp(x) + b exp(-x) y' = a exp(x) - b exp(-x) y'' = a exp(x) + b exp(-x) 1y = a exp(x) + b exp(-x) -4y' = -4a exp(x) +4b exp(-x) 3y'' = 3a exp(x) + 3b exp(-x) --------------------------- e(-x)+e(x) = 0a e(x) + 8b e(-x) b = 1/8 but that 0 has me a little iffied

OpenStudy (anonymous):

ok it seems correct

OpenStudy (anonymous):

thanks for your help

OpenStudy (amistre64):

might have to consider an "ax" instead of an a y = ax e(x) + b e(-x) y' = a e(x) + ax e(x) - b e(-x) y'' = 2a e(x) + ax e(x) + b e(-x) y = ax e(x) + b e(-x) -4y' = -4a e(x) -4ax e(x) +4b e(-x) +3y'' = 6a e(x) + 3x e(x) + 3b e(-x) --------------------------- = 2a e(x) + 0 +8b e(-x) b = 1/8\ a = -1/2

OpenStudy (anonymous):

yes this is what i found too. Thanks

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

but I remain unconvinced of the form of y

OpenStudy (amistre64):

y = yh + yp

OpenStudy (amistre64):

if you know how to play with a wronskian setup Wx W Wy e(3x) 0 e(x) 3e(3x) (e(x)+e(-x)) e(x) Wx = -(e(2x) + 1) Wy = (e(4x)+e(2x)) W = -2e(4x)